OpenStudy (anonymous):

solve the system equations. x+9y=2 (1) x=9-9y(2)

OpenStudy (anonymous):

x+9y=2 -x-9y=-9 these lines are parallell 0=-7 controdicts it self which means they never cross

OpenStudy (anonymous):

there are an infinite number of solutions to these equations

OpenStudy (anonymous):

y=-x/9+2/9 does not equal y=-x/9+1 when x=0 there are two different y values so they do not cross since they have the same slope